The Ratio of men:women:children living in Outapi is 6:7:3. There are 42 000 women
1.1 how many people altogether live in Outapi?
1.2 How many Children Live in Outapi?

Answer by Theo(13342)

the ratio of men to women to children is 6:7:3.

this means that, out of a total of every 16 people, 6 are men, 7 are women, 3 are children.

there are 42,000 women.

if the ratio of men to women is 6:7, this means that 6/7 = x/42,000.

x represents the number of men in this ratio.

cross multiply to get 6 * 42,000 = 7 * x

solve for x to get x = 6 * 42,000 / 7 = 36,000.

if the ratio of women to children is 7:3, this means that 7/3 = 42,000 / x.

x represents the number of children in this ratio.

cross multiply to get 7 * x = 3 * 42,000

solve for x to get x = 3 * 42,000 / 7 = 18,000.

we now know that there are 36,000 men and 42,000 women and 18,000 children.

the ratio of men to women is 36,000 / 42,000 = 36 / 42 = 6 / 7.

the ratio of women to children is 42,000 / 18,000 = 42 / 18 = 7 / 3.

the ratio of men to children is 36,000 / 18,000 = 36 / 18 = 6 / 3.

the total of people is 36,000 men and 42,000 women and 18,000 children = 96,000.

the ratio of 6:7:3 means that, out of every 16 people, 6 are men and 7 are women and 3 are children.

divide 96,000 by 16 to get 6000.

there 6000 sets of 16 people in a total of 96,000 people.

each of these sets has 6 men and 7 women and 3 chldren.

6,000 * 6 = 36,000 men.

6,000 * 7 = 42,000 women.

6,000 * 3 = 18,000 children.

the ratios hold.

another way to look at it is that there are 6 men for every 16 people, so the ratio of men to the total is 6/16, and the number of men total is 6/16 * 96,000 = 36,000.

likewise, there are 7 women for every 16 people, so the ratio of women to the total is 7/16, and the number of women total is 7/16 * 96,000 = 42,000.

likewise, there are 3 children for every 16 people, so the ratio of children to the total is 3/16, and the number of children total is 3/16 * 96,000 = 18,000.

any questions, email dtheophilis@gmail.com